A ball of mass 4 kg moves with a speed of 6 m/s. a) What is its momentum? b) What impulse is needed to stop it? c) What average force is needed to stop it in 3 s?
1
Expert's answer
2020-04-30T10:39:37-0400

(a) The momentum of a ball

"p_i=mv_i=(4\:\rm kg)\times (6\:\rm m/s)=24\: kg\cdot m/s"

(b) The impulse needed to stop  a ball


"\Delta p=|p_f-p_i|=|0-24|\:\rm kg\cdot m/s=24\: kg\cdot m/s"


(c) The average force needed to stop a ball

"F=\frac{\Delta p}{\Delta t}=\frac{24\:\rm kg\cdot m/s}{3\:\rm s}=8\:\rm N."